OpenStudy (anonymous):

Can someone please help me answer the questions: Find the exact value: (2- √2)² - (2+√2)² = ?

OpenStudy (anonymous):

a^2-b^2 = (a+b)(a-b) (2- √2)² - (2+√2)² =(2-(sqrt2)+2+(sqrt2))(2-(sqrt2)-2-(sqrt2)) =(4)(-2sqrt2) =-8sqrt2

OpenStudy (anonymous):

great :) thanks for replay!

OpenStudy (anonymous):

you are welcome